OpenStudy (anonymous):

Simplify: e^3ln(2x)+2ln(y^2)-4ln(2xy)

OpenStudy (anonymous):

does it help to know that \[e^{\ln(whatever)}=whatever\]?

OpenStudy (anonymous):

Sort of, but what do you do with the numbers before the ln? Just multiply them by whatever?

OpenStudy (sidsiddhartha):

2x^3-4ln(2x)

OpenStudy (shiraz14):

Am I correct to say that what you actually meant was the following: \[e ^{3\ln(2x) + 2\ln(y^{2}) - 4\ln(2xy)}\]

OpenStudy (anonymous):

Yes, that is what I meant.

OpenStudy (shiraz14):

(8x^3y^4) / (16x^4y^4) = 1/2x

OpenStudy (sidsiddhartha):

then its 1/2x

OpenStudy (anonymous):

Thank you both